OpenStudy (anonymous):

What is the first step in solving the equation 4x + 16 = 20? Divide both sides of the equation by −4. Divide both sides of the equation by −16. Subtract 16 from both sides of the equation. Add 16 to both sides of the equation.

OpenStudy (anonymous):

SUTRACT 16 FROM BOTH SIDES

OpenStudy (anonymous):

MEDAL PLZ IF I HELPED YOU

OpenStudy (anonymous):

4X+16-16=20-16 4X=4 X=1 THAT'S IT

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

AND MEDAL

OpenStudy (anonymous):

CLICK ON BEST RESPONSE BELOW BLUE COLORED BOX

OpenStudy (anonymous):

.

OpenStudy (anonymous):

ON YOUR RIGHT SIDE WHERE MY REPLIES ARE AT EVERY REPLY THERE IS A BOX BEST RESPONSE

OpenStudy (anonymous):

THERE IS ALSO BLUE COLOURED BEST RESPONSE

OpenStudy (anonymous):

Did I do it right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thanks